# Neutrino mixing parameters and masses from $\Delta(96)\rtimes H_{CP}$ in the tri-direct CP approach

### v1

@@ Abstract
We present a comprehensive model independent analysis of all breaking patterns resulting from $\Delta(96)\rtimes H_{CP}$ in the tri-direct CP approach of the minimal seesaw model with two right-handed neutrinos. The three generations of left-handed lepton doublets are assumed to transform as the irreducible triplet $\bm{3_{0}}$ of $\Delta(96)$, and the two right-handed neutrinos are assigned to singlets. In the case that both flavon fields $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ transform as triplet $\bm{\bar{3}_{0}}$, only one phenomenologically viable lepton mixing pattern is obtained for normal ordering neutrino masses. The lepton mixing matrix is predicted to be TM1 pattern, with neutrino masses, mixing angles, and CP violation phases depending on only three real input parameters. When $\phi_{\text{sol}}$ is assigned to the $\bm{\bar{3}_{1}}$ representation, an additional real parameter $x$ must be included. Then we find 42 independent phenomenologically interesting mixing patterns, and the corresponding predictions for lepton mixing parameters and neutrino masses are obtained. Furthermore, we perform a detailed numerical analysis for five example breaking patterns with some benchmark values of $x$. The absolute values of the first columns of PMNS matrix are fixed to be $\left(\sqrt{\frac{2}{3}},\frac{1}{\sqrt{6}},\frac{1}{\sqrt{6}}\right)^{T}$, $\frac{1}{5}\left(\sqrt{17},2,2\right)^{T}$, $\frac{1}{\sqrt{38}}\left(5,2,3\right)^{T}$, $\frac{1}{\sqrt{57}}\left(\sqrt{37},\sqrt{10},\sqrt{10}\right)^{T}$ and $\frac{1}{3}\left(\sqrt{6},1,\sqrt{2}\right)^{T}$, respectively.
@@ 1 Introduction
The confirmation of neutrino oscillations demonstrates that neutrinos possess tiny masses and that flavor mixing occurs within the lepton sector. The Standard Model (SM) does not provide any insight into the origin of fermion mass hierarchy, flavor mixing and CP violation. This remains a fundamental and unsolved flavor puzzle in physics. A definitive theoretical framework capable of resolving this puzzle has yet to emerge [1, 2, 3, 4, 5, 6]. The observation of non-zero neutrino masses, and large solar and atmospheric neutrino mixing angles open up a window to the new physics beyond SM. The discrete flavor symmetry approach, as a promising framework for the lepton flavor mixing angles, is now widely used to account for large neutrino mixing angles [7, 8, 9, 1, 2, 3, 10, 5, 4, 6, 11]. Non-Abelian discrete flavor symmetry combined with CP (gCP) symmetry provide a powerful framework to explain the lepton mixing angles and predict lepton CP violation phases [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. Furthermore, a simultaneous description of quark and lepton flavor mixing and CP violation can be achieved through spontaneous breaking of a discrete family symmetry and gCP symmetry [39, 40, 41]. The type I seesaw mechanism, which introduces heavy right-handed (RH) Majorana neutrinos, appears as the most elegant solution for generating tiny neutrino masses [45, 46, 47].
@@
Although the type I seesaw mechanism offers a qualitative explanation for the small neutrino mass scale via heavy right-handed neutrinos (RHNs). However, without additional assumptions, the three RHNs seesaw model contains too many parameters to yield specific predictions for neutrino mass and mixing. Furthermore, there is a mass hierarchy between the two mass squared differences $\Delta m^{2}_{21}\equiv m^{2}_{2}-m^{2}_{1}$ and $|\Delta m^{2}_{3l}|\equiv|m^{2}_{3}-m^{2}_{l}|$($\Delta m^{2}_{3l}\equiv m^{2}_{3}-m^{2}_{1}>0$ for normal ordering (NO) and $\Delta m^{2}_{3l}\equiv m^{2}_{3}-m^{2}_{2}<0$ for inverted ordering (IO)). To address these limitations, the sequential dominance (SD) framework [48, 49] provides an effective approach. It naturally results in the observed hierarchy between solar and atmospheric neutrino mass splittings $\Delta m^{2}_{21}$ and $\Delta m^{2}_{3l}$ [50] by imposing a strong mass hierarchy among the heavy Majorana neutrinos: $M_{\text{atm}}\ll M_{\text{sol}}\ll M_{\text{dec}}$. This originates from the proposal that a dominant heavy RHN generates the atmospheric neutrino mass, a heavier subdominant RHN generates the solar neutrino mass, and a potential third, largely decoupled RHN generates the lightest neutrino mass. This results in an effective two right-handed neutrino (2RHN) model with a massless lightest neutrino [51, 52], yielding simple and falsifiable predictions. To further increase predictive power of the minimal seesaw framework, models with one [53] or two [52] texture zeros in the Yukawa coupling were proposed. However, current data excludes two-zero models for normal ordering neutrino masses [54, 55, 56].
@@
The so-called CSD($n$) model [57, 58, 59, 60, 61, 62, 63, 64, 65, 66] represents a particularly predictive minimal seesaw model with one texture zero. Its Dirac neutrino mass matrix has two columns aligned with the directions $(0,1,-1)$ and $(1,n,2-n)$ respectively in the RHN diagonal basis, where the parameter $n$ was usually assumed to be a positive integer, but can generally be any real number. The CSD($n$) framework, including models like the CSD($3$) (also called Littlest Seesaw model) [59, 60, 61, 62, 63], CSD($4$) models [64, 65] and CSD($-1/2$[67] can give rise to phenomenologically viable predictions for lepton mixing parameters and neutrino masses as special cases of TM1 mixing in which the first column of the tri-bimaximal mixing matrix is preserved.
@@
In recent years, the idea of modular invariance [68] has provided a powerful framework to explain the lepton masses and mixing parameters, see Refs. [69, 70] for recent review. As was observed, modular symmetry indicates CSD($1-\sqrt{6}$) $\approx$ CSD($-1.45$) which requires multiple moduli [71, 72]. The first complete Littlest Modular Seesaw model based on CSD($1-\sqrt{6}$) was recently developed, where the three required moduli have been incorporated into complete models of leptons at the field theory level [73], or in 10-dimensional orbifolds [74]. A new related possibility based on CSD($1+\sqrt{6}$) $\approx$ CSD($3.45$) has been proposed [73]. Furthermore, the Littlest Modular Seesaw model has been extended to a Grand Unified scenario based on $SU(5)$ endowed with three modular $S_{4}$ symmetries [75].
@@
Based on the non-Abelian discrete flavor symmetry $G_{f}$ and gCP symmetry $H_{CP}$ compatible with it, we propose a new method called the tri-direct CP approach [76] for predicting neutrino masses and lepton mixing parameters in the minimal seesaw model with two right-handed neutrinos. In the tri-direct CP approach, the high energy symmetry $G_{f}\rtimes H_{CP}$ is spontaneously broken down to $G_{l}$, $G_{\text{atm}}\rtimes H^{\text{atm}}_{CP}$ and $G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}$ in the charged lepton, “atmospheric” and “solar” right-handed neutrino sectors, respectively [76]. Here $G_{l}$ is assumed to be an Abelian subgroup of $G_{f}$ and it is capable of distinguishing the three generations. The residual subgroups $G_{\text{atm}}\rtimes H^{\text{atm}}_{CP}$ and $G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}$ fix the two column of Dirac neutrino mass matrix. In the framework of the tri-direct CP approach, all phenomenologically interesting breaking patterns which can arise from $S_{4}\rtimes H_{CP}$ have been discussed for both NO and IO neutrino mass spectrum [77]. The model construction along the tri-direct CP approach was also illustrated [76, 77, 67]. Furthermore, the phenomenologically viable golden Littlest Seesaw model with GR1 mixing which preserves the first column of the golden ratio mixing matrix has been obtained from the breaking of $A_{5}$.
@@
In the present work, we perform a comprehensive model independent analysis of lepton mixing patterns resulting from the breaking of $\Delta(96)\rtimes H_{CP}$ via the tri-direct CP approach, considering the NO case. The three generations of the left-handed (LH) lepton doublets are assumed to transform as the irreducible triplet $\bm{3_{0}}$ of $\Delta(96)$, and the 2RHN are assigned to the singlets $\bm{1_{m}}$ ($m=0,1$). Then flavon fields $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ are necessary to break the high energy symmetry $\Delta(96)\rtimes H_{CP}$ down to $G_{\text{atm}}\rtimes H^{\text{atm}}_{CP}$ and $G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}$ in the atmospheric and solar neutrino sectors, respectively. Furthermore, they must transform as $\bm{\bar{3}_{m}}$ to preserve $\Delta(96)$ and gCP symmetry. If both $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ are assigned to the triplet $\bm{\bar{3}_{0}}$, we find that only one independent lepton breaking pattern can accommodate the experimental data in lepton sector, in which lepton mixing parameters and neutrino masses depend on three real input parameters $|m_{a}|$, $r$ and $\eta$. As a consequence, the corresponding lepton mixing matrix is predicted to be TM1 pattern. When $\phi_{\text{sol}}$ transforms as $\bm{\bar{3}_{1}}$, one more real free parameter $x$ would be imposed. Our analysis reveals that 42 independent breaking patterns are consistent with experimental data on neutrino masses and mixing parameters. We subsequently conduct an detailed analysis of five phenomenologically interesting mixing patterns with benchmark values for parameters $x$ and $\eta$. We identify four distinct mixing matrices that differ from the TM1 mixing scheme.
@@
We emphasise that there are good motivations to study $\Delta(96)$ flavor symmetry. As a flavor symmetry, $\Delta(96)$ is well established and widely discussed [78, 79, 18, 80], and breaking it can produce well-known mixing matrices such as tri-bimaximal, bimaximal, and Toorop-Feruglio-Hagedorn (TFH). $\Delta(96)$ is a group with complex triplet and sextet representations, thus offers new model building possibilities absent in $A_{4}$, $S_{4}$ and $A_{5}$. Such models readily extend to grand unified schemes like $\Delta(96)\times SU(5)$ [81]. Moreover, complex triplets forbid the trivial unit mass matrix which would destroy the mass hierarchies (unlike in $A_{4}$, $S_{4}$ and $A_{5}$ where it would be allowed). While analysis of these grand unified models lies beyond our scope, the formal results established here enable future exploration.
@@
The remaining parts of this paper are organized as follows. The framework of the tri-direct CP approach for 2RHN models, based on the discrete flavor group $G_{f}$ and gCP symmetry, is reviewed in section 2. In section 3, we perform a model independent analysis of all possible lepton mixing patterns achievable from $\Delta(96)\rtimes H_{CP}$, and the phenomenological predictions for the lepton mixing angles, CP violation phases and neutrino masses are obtained. Five example breaking patterns are presented and a thorough numerical analysis is performed to discuss their phenomenology in section 4. We summarize our results and give the conclusion in section 5. The group theory of $\Delta(96)$ is presented in appendix A, including the conjugacy classes, the irreducible representations, the Kronecker products, the Clebsch-Gordan (CG) coefficients, the Abelian subgroups and corresponding invariant vacuum expectation values (VEVs). The appendix B describes the diagonalization of a general subdiagonal neutrino mass matrix.
@@ 2 Framework
It is highly non-trivial to combine a family symmetry $G_{f}$ with the gCP symmetry together. In order to consistently combine flavor symmetry $G_{f}$ with gCP symmetry $H_{CP}$, the so-called consistency conditions have to be fulfilled [12, 13]:
@@
where $\rho_{\bm{r}}(g)$ represents the matrix of group element $g$ in the irreducible representation $\bm{r}$ of $G_{f}$, while $X_{\bm{r}}$ is the gCP transformation matrix. Crucially, the relation between $g$ and $g^{\prime}$ must hold identically across all irreducible representations $\bm{r}$. This equation reveals that $X_{\bm{r}}$ induces a mapping from group element $g$ to $g^{\prime}$, thereby establishing a connection between gCP transformations and automorphisms $u:G_{f}\rightarrow G_{f}$. Moreover, it was recently shown that the physically well-defined CP transformations have to be given by class-inverting automorphism of $G_{f}$ [82]. Notably, the composite transformation $\rho_{\bm{r}}(h)X_{\bm{r}}$ ($\forall h\in G_{f}$) corresponds to an automorphism formed by combining $u$ with an inner automorphism $\mu_{h}:g\rightarrow hgh^{-1}$ [83, 20]. Hence the mathematical structure of the full symmetry group comprising family symmetry $G_{f}$ and gCP symmetry is in general a semi-direct product. Consequently, the complete symmetry group is $G_{f}\rtimes H_{CP}$.
In the present work, we shall concentrate on the flavor symmetry $G_{f}=\Delta(96)$ which can be generated by four generators $a$, $b$, $c$ and $d$. The group theory of $\Delta(96)$, the representation matrices of generators, all Abelian subgroups and all the CG coefficients in our basis are reported in appendix A. Now we determine the explicit form of these gCP transformation matrices in our working basis. After performing a comprehensive study the automorphism group of $\Delta(96)$, we find a generic class-inverting automorphism $u$ of the $\Delta(96)$ group, and its actions on the generators $a$, $b$, $c$ and $d$ are as follows
@@
This indicates that the gCP transformation compatible with $\Delta(96)$ shares an identical structure with the family group transformation in our chosen basis, although they act on a field multiplet in different ways: under group elements $\varphi(x)\stackrel{{\scriptstyle g}}{{\longmapsto}}\rho_{\bf{r}}(g)\varphi(x)$, whereas under CP transformation, $\varphi(x)\stackrel{{\scriptstyle CP}}{{\longmapsto}}X_{\bf r}\varphi^{*}(x_{P})=\rho_{\bf r}(g)\varphi^{*}(x_{P})$, where $x_{P}=(t,-\vec{x})$. This leads to the consequence that all coupling constants become real number in a $\Delta(96)\rtimes H_{CP}$ model, owing to the reality of the CG coefficients demonstrated in table LABEL:tab:Delta96_CG.

### v2

@@ Abstract
We present a comprehensive model independent analysis of all breaking patterns resulting from $\Delta(96)\rtimes H_{CP}$ in the tri-direct CP approach of the minimal seesaw model with two right-handed neutrinos. The three generations of left-handed lepton doublets are assumed to transform as the irreducible triplet $\bm{3_{0}}$ of $\Delta(96)$, and the two right-handed neutrinos are assigned to singlets. In the case that both flavon fields $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ transform as triplet $\bm{\bar{3}_{0}}$, only one phenomenologically viable lepton mixing pattern is obtained for normal ordering neutrino masses. The lepton mixing matrix is predicted to be the TM1 pattern, with neutrino masses, mixing angles, and CP violation phases depending on only three real input parameters. When $\phi_{\text{sol}}$ is assigned to the $\bm{\bar{3}_{1}}$ representation, an additional real parameter $x$ must be included. Then we find 42 (12) independent phenomenologically interesting mixing patterns for normal (inverted) ordering neutrino masses, and the corresponding predictions for lepton mixing parameters and neutrino masses are obtained. Furthermore, we perform a detailed numerical analysis for five (one) example breaking patterns with some benchmark values of $x$ for normal (inverted) ordering. For the five normal examples, the absolute values of the first columns of the PMNS matrix are fixed to be $\left(\sqrt{\frac{2}{3}},\frac{1}{\sqrt{6}},\frac{1}{\sqrt{6}}\right)^{T}$, $\frac{1}{5}\left(\sqrt{17},2,2\right)^{T}$, $\frac{1}{\sqrt{38}}\left(5,2,3\right)^{T}$, $\frac{1}{\sqrt{57}}\left(\sqrt{37},\sqrt{10},\sqrt{10}\right)^{T}$ and $\frac{1}{3}\left(\sqrt{6},1,\sqrt{2}\right)^{T}$, respectively. For the inverted example, the absolute value of the third column of the PMNS matrix is $\frac{1}{2\sqrt{11}}\left(1,5,3\sqrt{2}\right)^{T}$.
@@ 1 Introduction
The confirmation of neutrino oscillations demonstrates that neutrinos possess tiny masses and that flavor mixing occurs within the lepton sector. The Standard Model (SM) does not provide any insight into the origin of fermion mass hierarchy, flavor mixing and CP violation. This remains a fundamental and unsolved flavor puzzle in physics. A definitive theoretical framework capable of resolving this puzzle has yet to emerge [1, 2, 3, 4, 5, 6]. The observation of non-zero neutrino masses, and large solar and atmospheric neutrino mixing angles open up a window to the new physics beyond SM. The discrete flavor symmetry approach, as a promising framework for the lepton flavor mixing angles, is now widely used to account for large neutrino mixing angles [7, 8, 9, 1, 2, 3, 10, 5, 4, 6, 11]. Non-Abelian discrete flavor symmetry combined with CP (gCP) symmetry provides a powerful framework to explain the lepton mixing angles and predict lepton CP violation phases [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. Furthermore, a simultaneous description of quark and lepton flavor mixing and CP violation can be achieved through spontaneous breaking of a discrete family symmetry and gCP symmetry [39, 40, 41]. The type I seesaw mechanism, which introduces heavy right-handed (RH) Majorana neutrinos, appears as the most elegant solution for generating tiny neutrino masses [45, 46, 47].
@@
Although the type I seesaw mechanism offers a qualitative explanation for the small neutrino mass scale via heavy right-handed neutrinos (RHNs), without additional assumptions, the three RHNs seesaw model contains too many parameters to yield specific predictions for neutrino mass and mixing. Furthermore, there is a mass hierarchy between the two mass squared differences $\Delta m^{2}_{21}\equiv m^{2}_{2}-m^{2}_{1}$ and $|\Delta m^{2}_{3l}|\equiv|m^{2}_{3}-m^{2}_{l}|$($\Delta m^{2}_{3l}\equiv m^{2}_{3}-m^{2}_{1}>0$ for normal ordering (NO) and $\Delta m^{2}_{3l}\equiv m^{2}_{3}-m^{2}_{2}<0$ for inverted ordering (IO)). To address these limitations, the sequential dominance (SD) framework [48, 49] provides an effective approach. It naturally results in the observed hierarchy between solar and atmospheric neutrino mass splittings $\Delta m^{2}_{21}$ and $\Delta m^{2}_{3l}$ [50] by imposing a strong mass hierarchy among the heavy Majorana neutrinos: $M_{\text{atm}}\ll M_{\text{sol}}\ll M_{\text{dec}}$. This originates from the proposal that a dominant heavy RHN generates the atmospheric neutrino mass, a heavier subdominant RHN generates the solar neutrino mass, and a potential third, largely decoupled RHN generates the lightest neutrino mass. This results in an effective two right-handed neutrino (2RHN) model with a massless lightest neutrino [51, 52], yielding simple and falsifiable predictions. To further increase predictive power of the minimal seesaw framework, models with one [53] or two [52] texture zeros in the Yukawa coupling were proposed. However, current data excludes two-zero models for normal ordering neutrino masses [54, 55, 56].
@@
The so-called CSD($n$) model [57, 58, 59, 60, 61, 62, 63, 64, 65, 66] represents a particularly predictive minimal seesaw model with one texture zero. Its Dirac neutrino mass matrix has two columns aligned with the directions (0,1,1)(0,1,-1) and (1,n,2n)(1,n,2-n) respectively in the RHN diagonal basis, where the parameter $n$ was usually assumed to be a positive integer, but can generally be any real number. The CSD($n$) framework, including models like the CSD($3$) (also called Littlest Seesaw model) [59, 60, 61, 62, 63], CSD($4$) models [64, 65] and CSD(1/2-1/2[67] can give rise to phenomenologically viable predictions for lepton mixing parameters and neutrino masses as special cases of TM1 mixing in which the first column of the tri-bimaximal mixing matrix is preserved.
@@
In recent years, the idea of modular invariance [68] has provided a powerful framework to explain the lepton masses and mixing parameters, see Refs. [69, 70] for recent review. As was observed, modular symmetry indicates CSD(161-$\sqrt${6}) $\approx$ CSD(1.45-1.45) which requires multiple moduli [71, 72]. The first complete Littlest Modular Seesaw model based on CSD(161-$\sqrt${6}) was recently developed, where the three required moduli have been incorporated into complete models of leptons at the field theory level [73], or in 10-dimensional orbifolds [74]. A new related possibility based on CSD($1+\sqrt{6}$) $\approx$ CSD($3.45$) has been proposed [73]. Furthermore, the Littlest Modular Seesaw model has been extended to a Grand Unified scenario based on $SU(5)$ endowed with three modular $S_{4}$ symmetries [75].
@@
Based on the non-Abelian discrete flavor symmetry $G_{f}$ and gCP symmetry $H_{CP}$ compatible with it, we propose a new method called the tri-direct CP approach [76] for predicting neutrino masses and lepton mixing parameters in the minimal seesaw model with two right-handed neutrinos. In the tri-direct CP approach, the high energy symmetry $G_{f}\rtimes H_{CP}$ is spontaneously broken down to $G_{l}$, $G_{\text{atm}}\rtimes H^{\text{atm}}_{CP}$ and $G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}$ in the charged lepton, “atmospheric” and “solar” right-handed neutrino sectors, respectively [76]. Here $G_{l}$ is assumed to be an Abelian subgroup of $G_{f}$ and it is capable of distinguishing the three generations. The residual subgroups $G_{\text{atm}}\rtimes H^{\text{atm}}_{CP}$ and $G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}$ fix the two columns of Dirac neutrino mass matrix. In the framework of the tri-direct CP approach, all phenomenologically interesting breaking patterns which can arise from $S_{4}\rtimes H_{CP}$ have been discussed for both NO and IO neutrino mass spectrum [77]. The model construction along the tri-direct CP approach was also illustrated [76, 77, 67]. Furthermore, the phenomenologically viable golden Littlest Seesaw model with GR1 mixing which preserves the first column of the golden ratio mixing matrix has been obtained from the breaking of $A_{5}$ [78].
@@
In the present work, we perform a comprehensive model independent analysis of lepton mixing patterns resulting from the breaking of $\Delta(96)\rtimes H_{CP}$ via the tri-direct CP approach, considering both NO and IO neutrino mass spectrums. The three generations of the left-handed (LH) lepton doublets are assumed to transform as the irreducible triplet $\bm{3_{0}}$ of $\Delta(96)$, and the 2RHN are assigned to the singlets $\bm{1_{m}}$ ($m=0,1$). Then flavon fields $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ are necessary to break the high energy symmetry $\Delta(96)\rtimes H_{CP}$ down to $G_{\text{atm}}\rtimes H^{\text{atm}}_{CP}$ and $G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}$ in the atmospheric and solar neutrino sectors, respectively. Furthermore, they must transform as $\bm{\bar{3}_{m}}$ to preserve $\Delta(96)$ and gCP symmetry. For the NO case, if both $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ are assigned to the triplet $\bm{\bar{3}_{0}}$, we find that only one independent lepton breaking pattern can accommodate the experimental data in lepton sector, in which lepton mixing parameters and neutrino masses depend on three real input parameters $|m_{a}|$, $r$ and $\eta$. As a consequence, the corresponding lepton mixing matrix is predicted to be TM1 pattern. When $\phi_{\text{sol}}$ transforms as $\bm{\bar{3}_{1}}$, one more real free parameter $x$ would be imposed. Our analysis reveals that 42 independent breaking patterns are consistent with experimental data on neutrino masses and mixing parameters. We subsequently conduct a detailed analysis of five phenomenologically interesting mixing patterns with benchmark values for parameters $x$ and $\eta$. We identify four distinct mixing matrices that differ from the TM1 mixing scheme. Similarly, for the IO case, we find 12 phenomenologically viable breaking patterns when $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ are assigned to the representations $\bm{\bar{3}_{0}}$ and $\bm{\bar{3}_{1}}$, respectively. We then perform a detailed analysis of one particularly interesting example among them.
@@
We emphasise that there are good motivations to study $\Delta(96)$ flavor symmetry. As a flavor symmetry, $\Delta(96)$ is well established and widely discussed [79, 80, 18, 81], and breaking it can produce well-known mixing matrices such as tri-bimaximal, bimaximal, and Toorop-Feruglio-Hagedorn (TFH). $\Delta(96)$ is a group with complex triplet and sextet representations, thus offering new model building possibilities absent in $A_{4}$, $S_{4}$ and $A_{5}$. Such models readily extend to grand unified schemes like $\Delta(96)\times SU(5)$ [82]. Moreover, complex triplets forbid the trivial unit mass matrix which would destroy the mass hierarchies (unlike in $A_{4}$, $S_{4}$ and $A_{5}$ where it would be allowed). While analysis of these grand unified models lies beyond our scope, the formal results established here enable future exploration.
@@
The remaining parts of this paper are organized as follows. The framework of the tri-direct CP approach for 2RHN models, based on the discrete flavor group $G_{f}$ and gCP symmetry, is reviewed in section 2. In section 3, we perform a model independent analysis of all possible lepton mixing patterns achievable from $\Delta(96)\rtimes H_{CP}$, and the phenomenological predictions for the lepton mixing angles, CP violation phases and neutrino masses are obtained. Six example breaking patterns are presented and a thorough numerical analysis is performed to discuss their phenomenology in section 4. We summarize our results and give the conclusion in section 5. The group theory of $\Delta(96)$ is presented in appendix A, including the conjugacy classes, the irreducible representations, the Kronecker products, the Clebsch-Gordan (CG) coefficients, the Abelian subgroups and corresponding invariant vacuum expectation values (VEVs). The appendix B describes the diagonalization of a general block-diagonal neutrino mass matrix.
@@ 2 Framework
It is highly non-trivial to combine a family symmetry $G_{f}$ with the gCP symmetry. In order to consistently combine flavor symmetry $G_{f}$ with gCP symmetry $H_{CP}$, the so-called consistency conditions have to be fulfilled [12, 13]:
@@
where $\rho_{\bm{r}}(g)$ represents the matrix of group element $g$ in the irreducible representation $\bm{r}$ of $G_{f}$, while $X_{\bm{r}}$ is the gCP transformation matrix. Crucially, the relation between $g$ and $g^{\prime}$ must hold identically across all irreducible representations $\bm{r}$. This equation reveals that $X_{\bm{r}}$ induces a mapping from group element $g$ to $g^{\prime}$, thereby establishing a connection between gCP transformations and automorphisms $u:G_{f}\rightarrow G_{f}$. Moreover, it was recently shown that the physically well-defined CP transformations have to be given by a class-inverting automorphism of $G_{f}$ [83]. Notably, the composite transformation $\rho_{\bm{r}}(h)X_{\bm{r}}$ ($\forall h\in G_{f}$) corresponds to an automorphism formed by combining $u$ with an inner automorphism $\mu_{h}:g\rightarrow hgh^{-1}$ [84, 20]. Hence the mathematical structure of the full symmetry group comprising family symmetry $G_{f}$ and gCP symmetry is in general a semi-direct product. Consequently, the complete symmetry group is $G_{f}\rtimes H_{CP}$.
In the present work, we shall concentrate on the flavor symmetry $G_{f}=\Delta(96)$ which can be generated by four generators $a$, $b$, $c$ and $d$. The group theory of $\Delta(96)$, the representation matrices of generators, all Abelian subgroups and all the CG coefficients in our basis are reported in appendix A. Now we determine the explicit form of these gCP transformation matrices in our working basis. After performing a comprehensive study of the automorphism group of $\Delta(96)$, we find a generic class-inverting automorphism $u$ of the $\Delta(96)$ group, and its actions on the generators $a$, $b$, $c$ and $d$ are as follows
@@
This indicates that the gCP transformation compatible with $\Delta(96)$ shares an identical structure with the family group transformation in our chosen basis, although they act on a field multiplet in different ways: under group elements $\varphi(x)\stackrel{{\scriptstyle g}}{{\longmapsto}}\rho_{\bf{r}}(g)\varphi(x)$, whereas under CP transformation, $\varphi(x)\stackrel{{\scriptstyle CP}}{{\longmapsto}}X_{\bf r}\varphi^{*}(x_{P})=\rho_{\bf r}(g)\varphi^{*}(x_{P})$, where $x_{P}=(t,-\vec{x})$. This leads to the consequence that all coupling constants become real numbers in a $\Delta(96)\rtimes H_{CP}$ model, owing to the reality of the CG coefficients demonstrated in table LABEL:tab:Delta96_CG.